OpenStudy (anonymous):

Will medal factor the trinomal x^2-17x+16 A:(x-2)(x-8) B:(x-16)(x-1) C:(x-16)(x+1) D:(x-4)(x+4) E: none

Directrix (directrix):

Can you think of two numbers that multiply to 16 and at the same time add to -17?

OpenStudy (anonymous):

umm not really...

Directrix (directrix):

@Alice2112 What is the answer to this: 16 times 1 equals ? You don't need the Quadratic Formula.

OpenStudy (anonymous):

16

Directrix (directrix):

Good. What is the product of -16 and -1 ? @Alice2112

OpenStudy (anonymous):

i domt know

Directrix (directrix):

A negative times a negative is positive. So, -16 times -1 equals +16. And, -16 + (-1) = -17 so we can begin to factor.

Directrix (directrix):

x^2 -17x + 16 = x^2 -16x -1x + 16 Note that the -17x has been rewritten as -16x - 1x

Directrix (directrix):

x^2 -16x -1x + 16 = x * (x - 16) -1* *(x - 16)

Directrix (directrix):

(x - 16) will be factored out. This is known as factoring by grouping.

Directrix (directrix):

= x * (x - 16) -1* *(x - 16) = (x - 16) [ x - 1] or (x - 16) ( x - 1) @Alice2112 which option is this?

OpenStudy (anonymous):

B

Directrix (directrix):

Correct. On multiple choice problems like this one, you could multiply out all the options and see which one produces the given equation but that would be a lot of work.

OpenStudy (anonymous):

thanks